PartnerObserving Log for 2007 Feb 15/16

Observer: Tom Herbst, Giacomo Beccari
Inst. Support: John Hill, Dave Thompson
TOs: Ed Lundin & David Gonzalez

Beautifully clear...a little windy with snow blowing around at the base of the dome We have installed a provisional baffle to reduce the scattering from the chamfer of L2. See the attached plot for the improvement.

(Note from JohnHill: This annular baffle behind L2 in combination with the 930 mm diameter baffle we added in front of the LBC-Blue hub appears to have been succesful in reducing the excess sky background by ~20% [by reducing the scattering inside the hub by ~50%]. Beware that these baffles will also change the large scale structure of the flat field, so these data will not be well flattened by earlier masterflats. Also note that the L2 baffle is not well centered. I think this miscentering is the source of the excess vignetting seen at the bottom of Chip 1. I plan to replace the L2 baffle with a better and more centered version on Monday or Tuesday -- I've just been to the UA bookstore to get a good supply of black poster board.)

18:30 - local time - Preparing for sky flats

19:00 - Resetting mount encoders to begin observing...

19:45 - difficulties converging active optics - likely very poor seeing...John estimates 4 arcsec. At 20:15 ambient = -7 C, glass = -3 C, steel = +2 C. Measurements on M87 confirm 4 arcsec seeing.

20:40 - terrible seeing - doing a sky background measurement to determine if provisional baffle has improved things. Issue with slow transfer from Archive to Repository.

21:30 - it's clear out there, but there is fine snow blowing around - cannot open the side vent doors. Doing photometric fields - seeing still extremely poor

22:30 - seeing still very poor- LBCFPIA does not converge...

22:45 - seems to be fewer stars out there. Lots of snow blowing around. Some is coming into the dome on level 5. Outside temperature has gone up ca. 1.5 degrees C in the last hour. Carrying on photometry and background checks at B, V.

23:30 - I think that there is thin stuff up there...hard to tell. Seeing still crappy. No snow accumulating in dome. Dave T. also thinks it may be a bit hazy...

23:50 - TCS computer has crashed - trying power cycle...Back up at 00:20...Now -4 deg C. ca. 3 deg up in 3 hrs.

00:40 - seeing now ~2.5 arcsec

1:00 - Executing M81 BV - seeing 1.8 arcsec, occasionally somewhat better - sky check looks ok.

1:30 - Standard PG 1047 for LBCFPIA and photometry

1:50 - Milne SN 2007jc - suitable for 1.5 arcsec plus seeing.

2:05 - Seeing 1.5-1.8 arcsec. Going to N5907 - 2:10 LBC has shut itself off. All systems switched to "Enabled". Restarting...Go at 2:25

2:40 - Some confusion - only one of five exposures executed - discrepancy in the OB file - five offsets provided, only one exposure requested (manually edited?)

3:00 - Seeing typ. 1.6-1.8

3:30 - Still clear out - definitely snow blowing around from the trees. Do LBCFPIA and continue on 5907...

4:30 - Seeing still 1.6 plus. Sky conditions outside unchanged

4:55 - OB execution error - telescope problem. Apparently the LSS (logging system) has died. David G. tries to re-start...This is not the problem anyway...Restarting LBCB @ 5:04..

5:25 - Weather continues fine. Doing photometric standard field PG 323. Seeing ~1.8-2.0 arcsec

5:27 - Only ~20-30 minutes until twilight - do another Milne SNe field - A4

5:45 - Monitoring sky level with SkyFlatTestG.ob - sky coming up slowly - move 15' north of Milne field for flats (airmass 1.2 in the west)

6:00 - Started SkyFlatG_5 with 30 sec exposures, guiding. After 5, rotate 180 and repeat

6:10 - Definitely brightening in east

6:14 - Starting R sky flats @ 20 sec per.

6:25 - Starting B sky flats @15 sec per...too much....reducing to 2 sec / frame

6:35 - It is pristine out there...

6:45 - closing up to do biases...great night except the awful seeing...
I Attachment Action Size Date Who Comment
Baffle_Improvement.gifgif Baffle_Improvement.gif manage 15 K 17 Feb 2007 - 04:05 TomHerbst Plot of improved scattered light with provisional baffle
PinholeImage.jpgjpg PinholeImage.jpg manage 138 K 17 Feb 2007 - 04:02 TomHerbst Pinhole image showing excess light from around primary
Topic revision: r8 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback